OpenStudy (anonymous):

What is the root of P(x)=2x^3-6x^2-5x+15?

OpenStudy (anonymous):

I used a calculator. \[(x-3)(\sqrt{2}x+\sqrt{5})(\sqrt{2}x-\sqrt{5})\]

OpenStudy (anonymous):

yes me too and I got the same answer

OpenStudy (anonymous):

You can choose between -5, 0, 1, 3

OpenStudy (anonymous):

I would go with 5 or 3

OpenStudy (anonymous):

Since the roots are where the function crosses the x axis, the answer is 3.